# Nudging Nutrition: Lessons from the Danish “Fat Tax”

### v2

@@ Abstract
In October 2011, Denmark introduced the world’s first and, to date, only tax targeting saturated fat. However, this tax was subsequently abolished in January 2013. Leveraging exogenous variation from untaxed Northern-German consumers, we employ a difference-in-differences approach to estimate the causal effects of both the implementation and repeal of the tax on consumption and expenditure behavior across eight product categories targeted by the tax. Our findings reveal significant heterogeneity in the tax’s impact across these products. During the taxed period, there was a notable decline in consumption of bacon, liver sausage, salami, and cheese, particularly among low-income households. In contrast, expenditure on butter, cream, margarine, and sour cream increased as prices rose. Interestingly, we do not observe any difference in expenditure increases between high and low-income households, suggesting that the latter were disproportionately affected by the tax. After the repeal of the tax, we do not observe any significant decline in consumption. On the contrary, there was an overall increase in consumption for certain products, prompting concerns about unintended consequences resulting from the brief implementation of the tax.
@@
We express our gratitude for the financial support provided by the Independent Research Fund Denmark, grant 6109-00075B, and Helsefonden, grant 18-B-0130. Additionally, we acknowledge the valuable feedback received from Phillip Heiler, Toru Kitagawa, Michael Lechner, Volha Lazuka, Pedro Sant’Anna, Jeff Smith, and the participants of the 2021 DGPE workshop, as well as the participants of the “Applied Econometrics of Program Evaluation” course. Special thanks are extended to Joachim Dahl Steensbjerre for his exceptional research assistance.
@@ 1 Introduction
Cardiovascular diseases are the world’s primary cause of mortality, and their prevalence due to modifiable risk factors like unhealthy diets is rising (Roth et al., , 2020). Since reducing saturated fat intake has been linked to a lower risk of cardiovascular diseases (Hooper et al., , 2020), it is no surprise that the World Health Organization strongly recommended to keep saturated fat content below 10% of the total energy intake (World Health Organization, , 2023). In Denmark, the dietary intake of saturated fats equaled 15% in 2010, surpassing the recommended levels and exhibiting relatively higher values compared to its European counterparts (Eilander et al., , 2015). Specifically, meat was found to account for 19% of the total saturated fat intake, while dairy products contributed with 30%, and fat-based products such as butter and margarine contributed with 32%.
@@
The excessive intake of saturated fat was the main reason for introducing a tax on saturated fat, the so-called “fat tax”, which was in place in Denmark between October 2011 and January 2013. It was a national initiative, applying to foods with saturated fat content exceeding 2.3%, including fats like margarine and butter, meats, and dairy products, but excluding drinking milk and milk-based yogurts. Despite its significance, there is limited empirical evidence on the effectiveness of the Danish fat tax to shape consumer responses, leaving policymakers with little guidance.
The economic rationale behind the implementation of food taxes, which aim at improving the healthiness of diets, is to address the externalities associated with the consumption of unhealthy foods. The effectiveness of a tax hence depends on consumers actually changing behavior in accordance with the tax (Cawley et al., , 2019). Assessing the efficacy of food taxes remains a complex task, influenced by a range of factors that shape consumer behavior and producer responses. Moreover, the overall impact on public health is contingent on consumer decisions regarding available substitutes and the healthiness of these alternatives. A few studies have estimated the effect of fat taxes using food price elasticities to simulate the effect of a fat tax on consumer behavior (Chouinard et al., , 2007; Allais et al., , 2010; Nordström and Thunström, , 2009), but their findings lack a causal interpretation (Powell et al., , 2013). Furthermore, these analyses assume that the tax is fully passed on to the consumers, which might not match reality. Theoretical studies on the consumer responses to food taxes (Lusk and Schroeter, , 2012; Yaniv et al., , 2009; Gortmaker et al., , 2011) stress the need for more robust empirical evidence using real-world data. A limited number of studies evaluated the Danish fat tax using real-world data (Jensen and Smed, , 2013; Jensen et al., , 2016; Bødker et al., 2015a, ; Smed et al., , 2016), resulting in inconclusive findings. In a recent review, Lhachimi et al., (2020) point out that two of these studies (Jensen and Smed, , 2013; Jensen et al., , 2016) lack an analysis of the effects of the tax on household expenditure. More importantly, all of these studies lack a control group that is not exposed to the tax and therefore lack a causal interpretation of the effect of the tax on consumer behavior. This forms a notable gap that is addressed in this study.
Our study design is unique in that we have access to a Northern-German households panel, which we use to construct a control group from a geographic area that is culturally close to Denmark in terms of food consumption behaviors, but is not affected by the tax. This allows us to estimate the causal effect of the tax on consumption and expenditure behaviors for different food groups employing the difference-in-differences (DID) estimator proposed by Callaway and Sant’Anna, (2021). We use Callaway and Sant’Anna, (2021) as our estimator for three reasons: it allows us to control for observed household characteristics in a flexible way; it is doubly robust, and it provides conservative confidence intervals.
We contribute to the existing literature in several ways. First, having access to a well-defined control group adds a causal perspective to the limited literature evaluating the impact of the fat tax on consumption of foods high in saturated fat. Second, we can estimate the causal effect of the tax on expenditure behaviors and the prices actually paid by the consumer. Third, our unique data allows us to study how the effect of the tax evolves over time. In particular, we offer new insights into the aftermath of the abolishment of the tax in January 2013, shedding light on potential habit formation and unintended effects on prices in the post-tax period. Moreover, we are able to estimate the causal effect of the tax on prices for each product group and measure how the tax was actually passed over to consumers. Finally, we estimate effect heterogeneity with respect to income levels to see whether low-income households were more affected by the tax than higher-income households. This is crucial to gain insight into the regressivity of the tax.
Our identification strategy relies on the assumption that purchases in Denmark and Germany would have followed the same trends had the tax not been introduced. Even though this assumption is essentially untestable, we study its plausibility through various pre-trend tests. A second crucial assumption is that there are no anticipation effects. The tax was announced in March 2011 before it was implemented in October 2011. We do not find any empirical evidence of hoarding behaviors in the two quarters prior to the tax implementation. However, to completely exclude the presence of potential anticipation effects, such as hoarding, observations six months before the introduction of the tax are excluded. Furthermore, to strengthen the validity of our results, we perform robustness checks with respect to cross-border shopping by excluding Danish households living close to the border with Germany from the analysis. We further show that our results do not rely on controlling for observed covariates, as they remain virtually unchanged when we use an unconditional DID estimator.
Our main findings suggest that the tax was effective in reducing the consumption of only three of the products we studied: cheese, liver sausage, and salami. This reduction in consumption is only present during the tax period, and we do not find any persistent effects on these products after the tax was abolished. For butter, cream, margarine, and sour cream, we find a significant increase in expenditure, which does not translate into a reduction in consumption. Interestingly, these are the only products for which we find a significant increase in price during the tax period. This price increase persists in the post-tax period for butter and sour cream. For butter, margarine, and sour cream, we observe an overall increase in consumption in the post-tax period. Overall, our findings suggest that the tax has largely failed to induce a long-run reduction in consumption for most products. For butter, cream, margarine and sour cream, the tax actually induced an increase in consumption, at least in the two years following the tax’s abolishment that we are able to observe in our data.
Examining effect heterogeneity between high and low-income households reveals a noteworthy reduction in both consumption and expenditure on cheese and liver sausage, but this reduction is significant only for the latter product group. Furthermore, a significant decrease in bacon consumption is identified exclusively among low-income households. Notably, both high and low-income households exhibited a comparable increase in expenditure on butter and margarine. Finally, the post tax increase, which we find for butter, cream, margarine and sour cream, is largely driven by high-income households.
@@
The remainder of this paper is organized as follows: First, we discuss the institutional background of the Danish fat tax in Section 2. Subsequently, we offer an overview of the data, including details of our identification strategy and its underlying assumptions in Section 3. The results are presented in Section 4, with a focus on each product category individually, along with potential price mechanisms. Section 5 addresses concerns related to our identification strategy, while Section 6 explores potential heterogeneity in tax responses. The paper concludes in Section 7.
@@ 2 Background on the Danish fat tax
In October 2011, Denmark introduced a tax targeting saturated fat, with the aim of improving dietary health and with the expectation of raising a tax revenue of 1 billion DKK (156.5 million USD [^3] A conversion rate of 6.39 DKK per USD is used throughout the paper.) (Bødker et al., 2015a, ; Smed, , 2012). The tax amounted to 16 DKK (2.51 USD) per kilo of saturated fat on foods that exceeded a threshold limit of 2.3% saturated fat. This amounts to 20 DKK (3.13 USD) per kilo including the 25% value added tax. The threshold was set such that drinking milk and milk-based yogurts were excluded from the tax. The tax predominantly affected products such as meat, dairy, oil, margarine, butter, and butter blends as these products contain considerable amounts of saturated fat. Producers had the option to use standardized coefficients for meat products instead of the actual fat content in the products. These coefficients were animal-specific (that is, there was one coefficient for pork products, one for beef, etc.), and based on the average content of saturated fat in the various types of meat consumed in Denmark (Smed, , 2012). Generally these coefficients were used instead of specific rates for different meat cuts since the latter option had higher administration costs. The administration of the fat tax was burdensome and estimated to have a retail and whole-sale cost of 200 million DKK (approximately 31 million USD) according to a report by the Danish Chamber of Commerce and the Confederation of Danish Industry (Bødker et al., 2015b, ).
The Danish Tax Committee anticipated an increase in consumer prices up to 7% for meats, up to 33% for dairy, up to 35% for animal fat, up to 16% for cooking oils, up to 24% for margarine and up to 20% for blended butter products (Bødker et al., 2015a, ). Konkurrence- og Forbrugerstyrelsen, (2014) analyzed the price increase of 27 taxed products and found that only for six of these products the consumer prices increased by more than can be explained by increases in the general food price index. They observed an overshifting of the tax, i.e. retailers increased the price on meat by more than what could be justified by the tax, increasing their profit margin by 1% to 5%. This finding is supported by Jensen et al., (2016) who found a price increase of 13% to 16% for high-fat minced beef and cream products. These estimated increases were higher than the a priori expected tax-induced price increases. The same study found small negative effect on prices for lower-fat products (Jensen et al., , 2016).
@@
3 Data, identification strategy, estimation, and inference
@@ 3.1 Data
GfK maintains a representative panel with respect to geographic location, education, and age. When a household leaves the panel, it is replaced by a similar household with respect to the aforementioned characteristics. Participants in the panel are equipped with a hand scanning device, enabling them to document detailed information on their weekly grocery shopping. The diary keeper reports the volume of foods purchased for the household, the expenditure, date and time of the purchase, among other details. While surveys and self-reported data are often susceptible to recall bias, this is less likely to be a concern when using scanning devices. A scanning device can ease registration of purchase information, which subsequently might be experienced as less burdensome to the households than registering food consumption manually. Moreover, our data provides an advantage over store scanner data by incorporating purchases made abroad. However, we acknowledge that our data do not account for food acquired outside of stores, such as in restaurants. The purchase diaries are submitted to GfK weekly and thoroughly checked for correct completion and consistency. As highlighted by Smed, (2008), a comparison of the budget-shares of foods from the GfK consumer panel with budget-shares revealed from data of Statistics Denmark shows no major biases for the distribution of expenses between basic food categories, even though the home-scan panel generally under-report in terms of absolute values.
@@
In our analysis, we consider a wide variety of product categories targeted to a different degree by the tax, encompassing bacon, butter and butter blends (referred to as butter), hard cheeses (referred to as cheese), cream, liver sausage and paté (referred to as liver sausage), margarine, salami, and sour cream. The product categories were first and foremost chosen based on their substantial contribution to the saturated fat intake in Denmark. We aggregate different products into products categories in a way that ensures both comparability between Germany and Denmark as well as an adequate number of observations.
@@
Our main analysis uses data covering the period from January 2009 until December 2014. The data are aggregated to a quarterly household panel, encompassing 11 quarters before the tax introduction and 13 quarters after the tax introduction. Furthermore, we create product-specific panels where households that never purchase from the considered product category are excluded. If the number of zero purchases for a given product changes systematically over time differentially between Denmark and Northern Germany, our results could be affected by sample selection bias. However, Figure 9 below shows similar pre-trends in the average number of observations with zero outcome between Denmark and Northern Germany. In general, for all product, we do not find any systematic change in the difference in the number of zeros even in the post-treatment period. The fact that we use quarterly data reduces the number of excluded households and the potential of sample selection bias.
Figure 1: Bacon.
Figure 2: Butter.
Figure 3: Cheese.
Figure 4: Cream.
Figure 5: Liver sausage.
Figure 6: Margarine.
Figure 7: Salami.
Figure 1: Bacon.
Figure 2: Butter.
Figure 3: Cheese.
Figure 4: Cream.
Figure 5: Liver sausage.
Figure 6: Margarine.
Figure 7: Salami.
Figure 8: Sour cream.
Figure 9: The figures show the de-trended quarterly average number of observations with zero value in the outcome variable for Danish and Northern-German households, respectively.
@@
The data further contain background information on the participating households, which is based on an annual questionnaire filled in by the diary keeper. This includes information on the geographic region of residence, the age of the head of the household, the number of children below age 15, the household size, income level, gender of the diary keeper, and education level. The data set with German households contains similar information as the Danish data. We harmonize the variables that are reported differently in the two data set, such as income and education. [^6] The education level is based on the International Standard Classification of Education (ISCED). To harmonize the income variable, which is reported in different brackets in the two data sets, we create quintiles based on the midpoints of the income ranges in the data set that contains all products for Germany and Denmark. We create five different levels, representing very low income, low income, medium income, high income and very high income. For our analysis, we use the most recently observed pre-tax characteristics for a household. That is, if a household’s income is observed in both the years 2009 and 2010, we use the income level observed in 2010. Pre-treatment characteristics of the households in our study are summarized in Table 6 in Appendix A.2.777Smed et al., (2016) use the same data and compare the descriptive statistics to official numbers from Statistics Denmark. They observe that the panel contains relatively more households in the capital and households with a lower level of educational attainment. There are some notable differences between the Northern-German and Danish households. For instance, the age of the head of the household has a tendency to be higher in Germany compared to Denmark. Additionally, there is a noticeable gender distribution distinction among diary keepers, with 78% being female in Denmark as opposed to only 59% in Germany. Moreover, households in Germany tend to be larger compared to those in Denmark. Since the household characteristics listed in the table are included as control variables in our difference-in-differences strategy, we address the concern that differences in consumption trends between Germany and Denmark might be caused by imbalances in sample composition.
@@
To investigate consumers’ response to the introduction of the fat tax, we define three distinct outcomes of interest. Firstly, we define three variables to investigate consumption and expenditure behavior: (1) the average quarterly household consumption, measured by the average weight or volume in gr or ml purchased per quarter per household, (2) the average number of packages purchased per household, and (3) the average quarterly total expenditure per household in Euro cents. Note that we use product purchases as a proxy for consumption, recognizing that a purchase does not necessarily indicate consumption. To investigate the price changes caused by the tax, we construct a variable for the average quarterly price paid per 100 units (gr or ml depending on the product) in Euro cents, allowing us to investigate whether the tax was passed on to consumers by retailers and producers. [^8] Danish expenditure is converted from DKK to Euro cents using the rate 7.4405. Additionally, since quarters have different number of days we re-weight our outcomes such that each quarter consist of the same amount of days. Finally, we partial out the effect of seasoanl trends from all outcomes. To achieve this we simply run two separate linear regressions, one for each country, for each outcome on quarterly dummies and we use the residuals of these regressions as outcomes. [^9] The pre-tax and tax average quantities of our outcomes of interest are presented in Table 7 in Appendix A.3. Outliers are removed using Tukey’s fences with the constant $k=3$.
3.2 Identification strategy, estimation, and inference
The key strength of our study lies in the ability to observe a proper control group, enabling us to identify a causal effect of the Danish fat tax on consumption behavior. We observe comprehensive information on Northern-German households from Schleswig-Holstein and Hamburg. A suitable control group not only removes time trends, but also unobserved demand shocks (Cawley et al., , 2019). A notable advantage of a geographically close control group is the cultural proximity between the treated and control sample, increasing the likelihood that they are subject to similar external shocks. Our identification strategy relies on the DID framework. The main idea is to compare Danish households to their German counterparts, both before and after the introduction of the fat tax, in terms of our outcome variable and for the different product categories. DID identifies the average treatment effect on the treated (ATT), that is, the effect of the fat tax on the Danish consumers.
Another noteworthy advantage of our study stems from the availability of longitudinal data, which allows us to study how consumption behavior varies with the length of exposure to the tax and how the effects evolve over time, effectively embracing an event-type study approach. We adopt the estimator for event-study designs proposed by Callaway and Sant’Anna, (2021), which extends the doubly robust estimator of Sant’Anna and Zhao, (2020) to the setting with multiple time periods. [^10] Note that Callaway and Sant’Anna, (2021) additionally consider the setting with variation in treatment timing (staggered adoption), where individuals belong to a certain group depending on the timing of their treatment. However, in our context, the number of groups equals one as the tax was simultaneously introduced and repealed for all individuals. As we have already mentioned in the introduction, the main advantage of this approach over other estimators, such as the common two-way fixed effects estimator (see, for instance, Roth et al., (2023) for a recent review of the DID literature), is that it allows to control for observed covariates in a flexible way.

### v3

@@ Abstract
In October 2011, Denmark introduced the world’s first and, to date, only tax targeting saturated fat. However, this tax was subsequently abolished in January 2013. Leveraging exogenous variation from untaxed Northern-German consumers, we employ a difference-in-differences approach to estimate the causal effects of both the implementation and repeal of the tax on consumption and expenditure behavior across eight product categories targeted by the tax. Our findings reveal significant heterogeneity in the tax’s impact across these products. During the taxed period, there was a notable decline in consumption of bacon, liver sausage, and cheese, particularly among low-income households. In contrast, expenditure on butter, cream, and margarine increased as prices rose. Interestingly, we do not observe any difference in expenditure increases between high and low-income households, suggesting that the latter were disproportionately affected by the tax. After the repeal of the tax, we do not observe any significant decline in consumption. On the contrary, there was an overall increase in consumption for certain products, prompting concerns about unintended consequences resulting from the brief implementation of the tax. Finally, we find strong evidence on an overall increase purchases of butter abroad for households living less than 50 km from the German boarder but we do not find strong evidence of spatial heterogeneous effects of the tax.
@@
We express our gratitude for the financial support provided by the Independent Research Fund Denmark, grant 6109-00075B, and Helsefonden, grant 18-B-0130. Additionally, we acknowledge the valuable feedback received from Phillip Heiler, Volha Lazuka, Toru Kitagawa, Michael Lechner, Pedro Sant’ Anna, Jeff Smith, Anthony Wray, and the participants of the 2021 DGPE workshop, the participants of the “Applied Econometrics of Program Evaluation” course, as well as the participants of the 2024 EALE conference. Special thanks are extended to Kasper Veje Jakobsen, Lajza Prekazi, and Joachim Dahl Steensbjerre for their exceptional research assistance.
@@ 1 Introduction
Taxation has long been employed as a policy instrument to influence consumer behavior, correct market failures, and promote social welfare. In economics, taxes serve as incentives or disincentives, aiming to internalize externalities associated with the consumption or production of certain goods (Pigou, , 1920). In the context of public health, taxes on tobacco, alcohol, sugary beverages, and unhealthy foods are designed to reduce consumption of harmful products, thereby improving health outcomes and reducing associated healthcare costs (Chaloupka and Warner, , 2000; Allcott et al., , 2019).
Economic theory suggests that imposing a tax on goods with negative externalities should lead to a decrease in their consumption, as higher prices deter consumers (Pigouvian taxes). Additionally, such taxes can generate government revenue that can be allocated to further mitigate the externality. However, the effectiveness of these taxes depends on the price elasticity of demand for the taxed goods, the availability of substitutes, and the ability of consumers and producers to adjust their behavior (Chetty et al., , 2009).
Empirical studies have documented both intended and unintended consequences of taxation policies. For instance, taxes on sugar-sweetened beverages (SSBs) have been shown to reduce consumption (Falbe et al., , 2016; Fichera et al., , 2021), but also, in some contexts, to lead to substitution effects where consumers switch to other unhealthy products (Miracolo et al., , 2021). Cross-border shopping to evade taxes is another unintended consequence that can undermine the policy’s effectiveness (Beatty and Shimshack, , 2011). Moreover, the regressive nature of such taxes can disproportionately affect lower-income households, raising equity concerns (Allcott et al., , 2019; Wright et al., , 2017).
@@
The Danish “fat tax,” implemented between October 2011 and January 2013, provides a compelling case study to examine both the intended and unintended effects of taxation as an incentive. Aimed at reducing the intake of saturated fats linked to cardiovascular diseases (Hooper et al., , 2020), the tax was levied on foods with saturated fat content exceeding 2.3%, including products like butter, margarine, meats, and dairy products, excluding drinking milk and milk-based yogurts (see, Smed, (2012) for more institutional details). Despite its potential significance for public health, there is limited empirical evidence on the effectiveness of the Danish fat tax in altering consumer behavior, leaving policymakers with little guidance on its actual impact.
Previous studies on food taxes have often relied on simulation models using price elasticities to predict consumer responses (Chouinard et al., , 2007; Allais et al., , 2010; Nordström and Thunström, , 2009), but these lack causal interpretations and may not capture real-world complexities (Powell and Chaloupka, , 2009). Moreover, many analyses assume full pass-through of the tax to consumer prices, which may not hold in practice due to strategic pricing by firms (Cawley and Frisvold, , 2017). Empirical evidence using real-world data is scarce, and existing studies on the Danish fat tax lack proper control groups, hindering causal inference (Jensen and Smed, , 2013; Bødker et al., 2015a, ; Smed et al., , 2016; Jensen et al., , 2016).
Our study addresses this gap by leveraging a unique dataset that includes both Danish households affected by the fat tax and a control group of Northern German households. Using a difference-in-differences (DID) approach (Callaway and Sant’Anna, , 2021), we estimate the causal effects of the fat tax on consumption and expenditure behaviors across different food groups. This methodology allows us to isolate the impact of the tax from other contemporaneous factors, providing robust evidence on both intended and unintended consequences.
Our findings contribute to the broader literature on taxation as an incentive by highlighting the nuanced effects of the Danish fat tax. Specifically, we find that while the tax was effective in reducing the consumption of only two products—cheese and liver sausage—it did not significantly reduce consumption of other high-fat products like butter, cream, and margarine. Instead, for these products, we observe significant increases in expenditure without corresponding reductions in consumption. This suggests that the tax led to higher prices that consumers, especially those with lower price sensitivity, were willing to pay, indicating an inelastic demand for these goods.
Furthermore, our analysis reveals unintended consequences such as cross-border shopping and the regressive nature of the tax. We find substantial evidence of increased cross-border purchases of butter by Danish households living close to the German border during the tax period. Despite this increase in cross-border shopping, the overall consumption of butter did not decrease significantly, implying that these avoidance behaviors mitigated the intended effect of the tax on reducing saturated fat intake. Additionally, we observe geographical variations in the tax’s impact on expenditure, with households farther from the border experiencing higher price increases due to the tax being fully passed through to consumers in those areas.
Importantly, we also find that lower-income households were more affected by the tax, exhibiting significant reductions in consumption of certain products. This raises concerns about the equity implications of such policies, as the tax may disproportionately burden lower-income individuals.
By placing our results within the broader economic literature on taxation and consumer behavior, we shed light on the complexities of using taxes as incentives. Our study underscores the importance of considering demand elasticities, substitution effects, geographical factors, and heterogeneity among consumers when designing tax policies to achieve public health objectives without unintended adverse effects. Policymakers should account for the ease of cross-border shopping and the elasticity of demand in different regions to ensure the effectiveness and equity of such taxes.
@@
The remainder of this paper is organized as follows: In Section 2, we discuss the institutional background of the Danish fat tax. Subsequently, Section 3 offers an overview of the data, detailing our identification strategy and the underlying assumptions. The results are presented in Section 4, where we focus on each product category individually and examine potential price mechanisms. Section 5 addresses concerns related to our identification strategy, while Section6 explores potential heterogeneity in tax responses. In Section 7, we provide some descriptive evidence on cross-border shopping of taxed products, estimate geographical variations in the tax’s impact, and potential spillover effects on German households living close to the border. The paper concludes in Section 8.
@@ 2 Background on the Danish fat tax
In October 2011, Denmark introduced a tax targeting saturated fat, with the aim of improving dietary health and with the expectation of raising a tax revenue of 1 billion DKK (156.5 million USD [^3] A conversion rate of 6.39 DKK per USD is used throughout the paper.) (Bødker et al., 2015a, ; Smed, , 2012). The tax was 16 DKK (2.51 USD) per kilo of saturated fat on foods exceeding a threshold limit of 2.3% saturated fat. Including the 25% value-added tax imposed on all products in Denmark, this increases to 20 DKK (3.13 USD) per kilo. The threshold was set such that drinking milk and milk-based yogurts were excluded from the tax. The tax predominantly affected products such as meat, dairy, oil, margarine, butter, and butter blends as these products contain considerable amounts of saturated fat. Producers had the option to use standardized coefficients for meat products instead of the actual fat content in the products. These coefficients were animal-specific (that is, there was one coefficient for pork products, one for beef, etc.), and based on the average content of saturated fat in the various types of meat consumed in Denmark (Smed, , 2012). Generally these coefficients were used instead of specific rates for different meat cuts since the latter option set higher requirements for documentation and thereby had higher administration costs. The administration of the fat tax was burdensome and estimated to have a retail and whole-sale cost of 200 million DKK (approximately 31 million USD) according to a report by the Danish Chamber of Commerce and the Confederation of Danish Industry (Bødker et al., 2015b, ).
The Danish Tax Committee anticipated an increase in consumer prices up to 7% for meats, up to 33% for dairy, up to 35% for animal fat, up to 16% for cooking oils, up to 24% for margarine and up to 20% for blended butter products (Bødker et al., 2015a, ). Konkurrence- og Forbrugerstyrelsen, (2014) analyzed the price changes of 27 taxed products and found that, for only six of these products, the consumer price increases exceeded what could be explained by the general food price index trends. For some of these products, they observed an overshifting of the tax, i.e., retailers increased the price on meat by more than what could be justified by the tax, increasing their profit margin by 1% to 5%. This finding is supported by Jensen et al., (2016) who found a price increase of 13% to 16% for high-fat minced beef and cream products. These estimated increases were higher than the a priori expected tax-induced price increases. The same study found small negative effect on prices for lower-fat products (Jensen et al., , 2016).
@@
3 Identification strategy, data, estimation, and inference
3.1 Identification strategy
Our study’s key strength lies in observing a suitable control group—Northern German households from Schleswig-Holstein and Hamburg—which enables us to identify the causal effect of the Danish fat tax on consumption behavior. This control group helps remove time trends and unobserved demand shocks (Cawley et al., , 2019), and their geographical proximity and cultural similarity to Danish households increase the likelihood of experiencing similar external shocks. The regions of Schleswig and Holstein (Now Schleswig-Holstein in Germany) as well as Northern Schleswig (now Southern Jutland in Denmark) have been under alternating Danish and German governance since the Middle Ages. This intertwined governance led to a unique blend of Danish and German cultures in these areas, leaving Danish minorities in German territories and German minorities in Danish regions.
Today, the Danish-German border region is therefore characterized by significant cultural and economic integration. The Danish minority in South Schleswig, Germany, and the German minority in North Schleswig, Denmark, enjoy protections and support from both governments. The Copenhagen-Bonn Declarations of 1955 formalized minority rights, allowing for the preservation of language, education, and cultural practices. These minorities operate their own schools, libraries, churches, and cultural institutions, fostering strong cultural ties across the border. Economic connections are also robust, with cross-border trade and labor mobility contributing to the regional economy. Shared industries, agricultural practices, and retail markets further intertwine the regions economically.
These deep cultural and economic ties between Denmark and Northern Germany make the latter an appropriate control group for our study. The similarities in consumer behavior, dietary patterns, and market dynamics enhance the validity of our DID approach.
@@ 3.2 Data
GfK maintains a representative panel with respect to geographic location, education, and age.  [^6] See, https://panel.gfk.com/scan-dk/hjem When a household leaves the panel, it is replaced by a similar household with respect to the aforementioned characteristics. Participants in the panel are equipped with a hand scanning device, enabling them to document detailed information on their weekly grocery shopping. The diary keeper reports the volume of foods purchased for the household, the expenditure, date and time of the purchase, among other details. While surveys and self-reported data are often susceptible to recall bias, this is less likely to be a concern when using scanning devices. A scanning device can ease registration of purchase information, which subsequently might be experienced as less burdensome to the households than registering food consumption manually. Moreover, our data provides an advantage over store scanner data by incorporating purchases made abroad. However, we acknowledge that our data do not account for food acquired outside of stores, such as in restaurants. The purchase diaries are submitted to GfK weekly and thoroughly checked for correct completion and consistency. As highlighted by Smed, (2008), a comparison of the budget-shares of foods from the GfK consumer panel with budget-shares revealed from data of Statistics Denmark shows no major biases for the distribution of expenses between basic food categories, even though the home-scan panel generally under-report in terms of absolute values.
@@
In our analysis, we consider a wide variety of product categories targeted to a different degree by the tax, encompassing bacon, butter and butter blends (referred to as butter), hard cheeses (referred to as cheese), cream, liver sausage, margarine, salami, and sour cream.  [^7] We found no significant differences between butter and butter blends, so we combined them into a single category. The product categories were first and foremost chosen based on their substantial contribution to the saturated fat intake in Denmark, secondly based on comparability between Germany and Denmark as well as an adequate number of observations.
@@
Our main analysis uses data covering the period from January 2009 until December 2014. The data are aggregated to a quarterly household panel, encompassing 11 quarters before the tax introduction and 13 quarters after the tax introduction. Furthermore, we create product-specific panels where households that never purchase from the considered product category are excluded. If the number of zero purchases for a given product changes systematically over time differentially between Denmark and Northern Germany, our results could be affected by sample selection bias. However, Figure 1 below shows similar pre-trends in the average number of observations with zero outcome between Denmark and Northern Germany. In general, for all product, we do not find any systematic change in the difference in the number of zeros even in the post-treatment period. This may be due to our use of quarterly data, which significantly reduces the number of excluded households and the potential for sample selection bias.
(a) Bacon.
(b) Butter.
(c) Cheese.
(d) Cream.
(e) Liver sausage.
(f) Margarine.
(g) Salami.
(h) Sour cream.
Figure 1: Trends in the average number of zero-outcome observations across various products between Denmark and Northern Germany.
@@
The data further contain background information on the participating households, which is based on an annual questionnaire filled in by the diary keeper. This includes information on the geographic region of residence, the age of the head of the household, the number of children below age 15, the household size, income level, gender of the diary keeper, and education level. We harmonize the variables that are reported differently in the Danish and German data sets, such as e.g. income and education. [^8] The education level is based on the International Standard Classification of Education (ISCED). To control for income, we harmonize the income variable, as it is reported in different intervals in the two datasets. To ensure a valid comparison of households with similar income levels, we generate separate quintiles for Germany and Denmark using the midpoints of each income range, incorporating all available households in each country. These quintiles categorize households into five income levels: very low, low, medium, high, and very high income. For our analysis, we use the most recently observed pre-tax characteristics for a household. That is, if a household’s income is observed in both the years 2009 and 2010, we use the income level observed in 2010. Pre-treatment characteristics of the households in our study are summarized in Table LABEL:tab:sum_stat in Appendix A.2.999Smed et al., (2016) use the same data and compare the descriptive statistics to official numbers from Statistics Denmark. They observe that the panel contains relatively more households in the capital and households with a lower level of educational attainment. There are some notable differences between the Northern-German and Danish households. For instance, the head of the household tend to be older in the Germany data compared to the Danish data. Additionally, there is a noticeable gender distribution distinction among diary keepers, with 78% being female in Denmark as opposed to only 59% in Germany. Moreover, households in Germany tend to be larger compared to those in Denmark. Since the household characteristics listed in the table are included as control variables in our difference-in-differences strategy, we address the concern that differences in consumption trends between Germany and Denmark might be caused by imbalances in sample composition.
@@
To investigate consumers’ response to the introduction of the fat tax, we define three distinct outcomes of interest. Firstly, we define three variables to investigate consumption and expenditure behavior: (1) the average quarterly household consumption, measured by the average weight or volume in gr or ml purchased per quarter per household, (2) the average number of packages purchased per household, and (3) the average quarterly total expenditure per household in Euro cents. Note that we use product purchases as a proxy for consumption, recognizing that a purchase does not necessarily indicate consumption. To investigate the price changes caused by the tax, we construct a variable for the average quarterly price paid per 100 units (gr or ml depending on the product) in Euro cents, allowing us to investigate whether the tax was passed on to consumers by retailers and producers. [^10] Danish expenditure is converted from DKK to Euro cents using the rate 7.4405. Additionally, since quarters have different number of days we re-weight our outcomes such that each quarter consist of the same amount of days. Finally, we partial out the effect of seasonal trends from all outcomes. To achieve this we simply run two separate linear regressions, one for each country, for each outcome on quarterly dummies and we use the residuals of these regressions as outcomes. [^11] The pre-tax and tax average quantities of our outcomes of interest are presented in Table 9 in Appendix A.3. Outliers are removed using Tukey’s fences with the constant $k=3$.
3.3 Causal Effects Definition and Estimation